# Censored extreme value estimation

### v1

@@ Abstract
A novel and comprehensive methodology designed to tackle the challenges posed by extreme values in the context of random censorship is introduced. The main focus is the analysis of integrals based on the product-limit estimator of normalized top-order statistics, denoted extreme Kaplan–Meier integrals. These integrals allow for transparent derivation of various important asymptotic distributional properties, offering an alternative approach to conventional plug-in estimation methods. Notably, this methodology demonstrates robustness and wide applicability within the scope of max-domains of attraction. An additional noteworthy by-product is the extension of residual estimation of extremes to encompass all max-domains of attraction, which is of independent interest.
@@
Kaplan-Meier estimator,
@@ 1 Introduction
The foundational statistical work in fully-observed extreme value theory is by now classic and well-understood in the univariate case, see [2, 19, 15], and to a large degree for univariate time series, as can be found in [30]. Similarly, the pillars of survival analysis have been laid out since the seventies, in different flavors, with the counting process point of view being one of the the most popular approaches, and with martingale techniques playing an ubiquitous role, see [24, 1].
A staple of nonparametric estimation of the survival function for randomly censored time-to-event data is the Kaplan–Meier product-limit estimator [29], which has received continued attention through the decades. In [13] the rate of strong uniform consistency for the product-limit estimator was studied, and in [27] a law of the iterated logarithm was provided. The functional point of view has been explored in [16], providing functional limit laws for the increment functions of empirical processes based on randomly right-censored data. Concentrating on the tail empirical process of a fully-observed sample, the results of [10] provide the main tools to assert asymptotic functional distributional properties for extreme estimators. The empirical process approach has also resulted in fruitful limit theorems for time-series, as treated in [30].
Integral functions of the empirical distribution function were studied in [11] in terms of normal and stable limits. The generalization of this idea turned out to be paramount for survival analysis, integrating with respect to the Kaplan–Meier estimator. Indicator functions were first treated in [36] and [41], while the mean residual lifetime, corresponding to integrating the identity function, was treated in [40]. In the latter, the main mathematical technique was integration by parts. In [39] arbitrary integrals with respect to the Kaplan–Meier estimator were considered, establishing strong convergence through martingale techniques, and then [38] provided the central limit theorem using U-statistics. These results provide a general framework for dealing with distributional estimators from censored samples, and importantly, seamlessly deal with convergence on the largest possible support – the maximal interval – by appealing to U-statistics theory. Convergence of the product-limit estimator on the maximal interval was first dealt with successfully in [41], see also [42] for random truncation.
@@
Integral functions with respect to the empirical distribution function of top order statistics has also received interest. Such integrals can often be cast as sums of extreme values from regularly varying distributions, which was studied in [14] for regularly-varying tails and in [12] for arbitrary tail behavior. Upon replacement of the empirical distribution function with the Kaplan–Meier estimator of top-order statistics of a censored sample, the analysis of integral functions is no longer transparent with empirical processes theory, nor using the sums approach. This paper is devoted to studying precisely these integral functions by using the elegant U-statistics approach of [38], together with machinery from extreme value theory.
@@
More precisely, this paper proposes the analysis of randomly right-censored extremes through Kaplan–Meier integrals of normalized top-order statistics. In order to generalize the approach of [38] to the extreme case, the theory of U-statistics [35] and of residual estimation of [34] are of utmost importance. However, they are not directly applicable, the former having to be adapted to accommodate upper order statistics (see [32] for a related concept), and the latter requiring extension to the Gumbel and Weibull max-domains of attraction. The result provides a unifying and transparent treatment of extreme estimators and their convergence, imposing natural conditions. Note that extreme value index estimators derived from extreme Kaplan–Meier integrals have similar form and in some cases equal asymptotic variance as the ones treated in [6], though the construction principle is rather different and our theory allows for the construction of further estimators. In particular we provide a detailed account on a novel construction of the censored version of the moment estimator [17] of the extreme value index for all max-domains of attraction.
The main theoretical contributions of the paper can be summarized as follows:
@@
    Suitable conditions are established under which extreme Kaplan–Meier integrals are weakly consistent and asymptotically normal within the Fréchet max-domain of attraction, that is for distributions with regularly varying tails. Asymptotic normality is provided in several steps, with only the last one requiring a second order condition. The main unavoidable restriction is a second (plus epsilon) moment condition for the limit variance to exist, and in particular there are no requirements on the rate of uniform convergence of the asymptotic censoring proportion.
@@
    Residual estimation theory is then generalized to all max-domains of attraction, and in particular it is found that convergence rates vary across domains. There results are presented for fully-observed data and are of independent interest. The exposition is in terms of so-called candidate functions, which themselves must satisfy a second-order type condition.
@@
    The results on residual estimators is then used to extend the distributional convergence of extreme Kaplan–Meier-type integrals for arbitrary max-domains of attraction. Here, a regularly varying index of certain cdf’s plays a central role. Such index can be identified as the asymptotic non-censoring proportion.
@@
    Finally, we provide full asymptotics for a novel censored version of the moment estimator of [17], using all the main components of our theory. Some assumptions from the general theory are relaxable for this application and thus some more technical proofs are provided for precisely this purpose.
The cross-fertilization of extreme value theory and survival analysis has been explored before in various contexts. We provide an account of some of the most important contributions. In [3] the Pareto index estimation was considered for the first time under random right-censorship framework (see also [4, 18] covering all max-domains of attraction), which has since produced some generalizations, e.g. using covariates in [31], truncation in [25], trimming in [9], and bias-reduction in [5, 26]. Another approach, which neatly formulates the censoring problem with extremes as a path-wise problem, was proposed in [43] through constructions based on expected values. Subsequently, a partial result on the asymptotic distribution was recently provided in [6]. However, their method of proof is not generalizable to our setting. In terms of missing extreme data, the main reference is [44], and finally the machinery of extreme value theory has recently been imported to survival analysis in a different context, namely for cure-rate modelling in the case of insufficient follow-up, see [20, 21, 22].
Organisation of the paper. In Section 2 we introduce notation and results from extreme value theory and survival analysis relevant to our setup. Section 3 is devoted to the study of extreme Kaplan–Meier integrals when both event and censoring variables have distributions with regularly varying tails; Theorem 3.3 provides a general decomposition, which is exploited in Theorem 3.4 for consistency and in a series of asymptotic normality results culminating in Corollary 3.10. Residual estimators and their applications are then studied in Section 4 with no max-domain constraints and extensions of the results of the previous section are provided; we highlight Theorem 4.11 and Theorem 4.12 as our main extensions. Finally, Section 5 provides a prime application of the previous sections, analysing in detail an extreme value index estimator for censored data and all max-domains of attraction; we provide in Theorem 5.1 consistency and in Theorem 5.2 asymptotic normality. Proofs may be found in Appendices A, B and C.
@@ 2.1 Random censoring
Let $X_{1},\ldots,X_{n}$ be i.i.d. random variables with common cdf $F.$
Let $\varphi:\amsmathbb{R}\to\amsmathbb{R}$ be some measurable function which is $F$-square integrable, that is $\int\varphi^{2}\mathop{}\!\mathrm{d}F<\infty$. Then, an application of the central limit theorem (CLT) for $\{\varphi(X_{i})\}_{i=1}^{n}$ can be written as
$\sqrt{n}\int\varphi\,\mathop{}\!\mathrm{d}(F_{n}-F)\stackrel{{\scriptstyle d}}{{\to}}N(0,\sigma_{\varphi}^{2}),$ (2.1)
where $F_{n}$ is the empirical cdf of $\{X_{i}\}_{i=1}^{n}$ and $\sigma_{\varphi}^{2}=\int\varphi^{2}\mathop{}\!\mathrm{d}F-\left(\int\varphi\,\mathop{}\!\mathrm{d}F\right)^{2}.$ We now consider a random right-censoring assumption, that is, we assume that we only can observe
@@
where $Y_{1},\ldots,Y_{n}$ is an i.i.d. censoring sequence with common cdf $G$ independent of $\{X_{i}\}_{i=1}^{n}$. Notice that $\delta_{i}$ indicates whether $X_{i}$ has been observed or not. Let $H$ denote the cdf of $Z_{1}$, which by independence of $X_{1}$ and $Y_{1},$ is given by
@@
A consistent estimator of $F$, [41], is now provided under by the Kaplan–Meier product-limit estimator $\amsmathbb{F}_{n}$ instead of $F_{n}$ in (2.1), defined by
@@
More precisely, write $F\{a\}=F(a)-F(a-)$, such that $F\{a\}>0$ iff $a$ is an atom of $F$. Let $A$ be the set of all atoms of $H,$ possibly empty. Under $\int|\varphi|\mathop{}\!\mathrm{d}F<\infty,$ Stute and Wang [39] showed that
@@
where $(Z,\delta)\stackrel{{\scriptstyle d}}{{=}}(Z_{1},\delta_{1}).$
The following three functions are necessary to state a key asymptotic decomposition of $\int\varphi\,\mathop{}\!\mathrm{d}\amsmathbb{F}_{n}$, and are stated here for future reference.
@@ 2.2 Extreme value theory
We are interested in analyzing the extreme value behaviour of the datapoints $X_{1},\dots,X_{n}$, which is possible through the study of $Z_{i,n}$ and $\delta_{[i:n]}$.
Define for any cdf $F$ its survival function as $\bar{F},$ that is $\bar{F}=1-F,$ and its tail counterpart as
@@
which is itself a distribution function. Further, let $\gamma^{t}_{0},$ $\gamma^{t}_{1}$ and $\gamma^{t}_{2}$ are defined similarly to (2.5), (2.6) and (2.7), by replacing $H,H^{0}$ and $H^{1}$ with $H^{t},$ $H^{0,t}$ and $H^{1,t},$ respectively, where
@@
Finally, for future reference, define
$C^{t}(x)=\int_{1}^{\infty}\frac{G^{t}(\mathop{}\!\mathrm{d}y)}{(1-H^{t}(y))(1-G^{t}(y))}.$
First, we consider the case where the variables $X_{i}$ and censoring mechanism $Y_{i}$ are both heavy-tailed. In later sections this restriction is relaxed. More specifically, we assume now that $F$ and $G$ belong to the Fréchet max-domain of attraction, corresponding to a positive tail parameter. This implies that $F,$ $G$, and hence $H$, are regularly varying cdfs with respective tail indices $1/\gamma_{F},$ $1/\gamma_{G}$ and $1/\gamma_{H},$ where by (2.2) we obtain
@@
In this case, by Theorem 1.2.1, [15],
$1-F^{t}(x)=\frac{1-F(tx)}{1-F(t)}\to x^{-1/\gamma_{F}}\quad\text{ as }\quad t\to\infty,$ (2.12)
and similarly for $G$ and $H$. Note that the latter convergence is uniform in $x>1$ by Potter’s bounds (see, e.g., Proposition B.1.9, [15]).
The sequence of integers $k_{n}$ is an intermediate sequence if
Standard results of extreme value theory indicate that for an intermediate sequence $k=k_{n}$, the normalized upper order statistics $\{Z_{n-k+i,n}/Z_{n-k,n}\}_{i=1}^{k}$ given $Z_{n-k,n}=t$ are distributed as a vector of order statistics $\{V^{\ast}_{i,k}\}_{i=1}^{k}$ of a sample $\{V^{\ast}_{i}\}_{i=1}^{k}$ having cdf $H^{t}$ that converges to the Pareto distribution with index $1/\gamma_{H}$ as $t\to\infty$. However, to study the distributional convergence to the Pareto law with the correct target index $1/\gamma_{F}$ it is convenient to consider a product-limit construction, as follows (see also [18] for a moment construction).
@@
The extreme Kaplan–Meier estimator (EKM) is given by
@@
3 Extreme Kaplan–Meier Integrals
The asymptotic behaviour of extreme Kaplan–Meier integrals is studied in this section, defined as integrals with respect to the EKM estimator:
@@ 3 Extreme Kaplan–Meier Integrals
For technical reasons, such as applying the dominated convergence theorem, we assume that $\varphi$ is an almost everywhere (with respect to Lebesgue measure, written as a.e. for brevity) continuous function on $[1,\infty)$ and that $F,G,$ and therefore $H$ are eventually continuous.
Definition 2.
A continuous and eventually monotone function $\bar{\varphi}:[1,\infty)\to\amsmathbb{R}_{+}$ such that $|\varphi|\leq\bar{\varphi}$, is called an envelope of $\varphi.$
In the sequel, whenever we impose conditions on an envelope, we tacitly assume that the envelope exists. Otherwise, the condition is violated. The use of envelopes is technical in nature, to avoid requiring continuity and eventual monotonicity directly on $\varphi$. For a simple $\varphi$ function, the corresponding envelope is usually itself.
The following condition serves as a building block towards the general case, and is initially imposed, though later removed.
Temporary Condition 1.
The function $\varphi$ satisfies
$\varphi(x)=0\>\>\text{ for all }\>\>x>T\>\>\text{ and some }\>\>T<\infty.$ (3.2)
Important issues when studying EKM integrals arise from the fact that the “sample” is no longer independent. As a preliminary step, let us consider the behaviour of the EKM integral (3.1) conditioning on the denominator, that is given $Z_{n-k,n}=t.$ In this case, by analogy with Lemma 3.4.1, [15], the joint distribution of the vector of normalized order statistics $(Z_{n-k+1,n}/Z_{n-k,n},\ldots,Z_{n,n}/Z_{n-k,n})$ given $Z_{n-k,n}=t$ equals the joint distribution of order statistics $(V^{\ast}_{1,k},\ldots,V^{\ast}_{k,k})$ corresponding to i.i.d. random variables $\{V^{\ast}_{i}\}_{i=1}^{k}$ with cdf $H^{t}.$ By the Skorokhod representation theorem (e.g., p. 70 in [7]), on a rich enough probability space, it is possible to select $\{V^{\ast}_{i}\}_{i=1}^{k}$ in such a way that given $Z_{n-k,n}=t$ it holds $Z_{n-k+i,n}/Z_{n-k,n}=V_{i,k}^{\ast}$ almost surely, $i=1,\ldots,k.$ The following notation is useful in the sequel.
Definition 3.
For the random vectors
$\bar{Z}^{\ast}_{k}=(Z^{\ast}_{1,k},\dots,Z^{\ast}_{k,k})=(Z_{n-k+1,n}/Z_{n-k,n},\ldots,Z_{n,n}/Z_{n-k,n}),$
an array of functions $(f_{k,n})_{k,n\in\amsmathbb{N}}$, the baseline random variable $Z_{n-k,n}$, and a deterministic sequence $a_{1},a_{2},\ldots$ we say that $f_{k,n}(\bar{Z}^{\ast}_{k})$ is $o_{{P}}(a_{k})$ uniformly in large $t$ w.r.t. $Z_{n-k,n}$ if for every $\varepsilon>0$,
$\lim_{n/k,\,k\to\infty}\limsup\limits_{t\rightarrow\infty}{\amsmathbb{P}}(|f_{k,n}(\bar{Z}^{\ast}_{k})/a_{k}|>\varepsilon|\,Z_{n-k,n}=t)=0.$ (3.3)
We write in short (when no confusion arises) that $f_{k,n}(\bar{Z}^{\ast}_{k})=\overline{o}_{P}(a_{k})$. Analogously $f_{k,n}(\bar{Z}^{\ast}_{k})=\overline{O}_{P}(a_{k})$ using the notions of bounded in probability in place of convergence to zero in probability.
Note that $\bar{Z}^{\ast}_{k}\stackrel{{\scriptstyle d}}{{=}}\bar{V}^{\ast}_{k}$ given $Z_{n-k,n}=t.$ In the proofs, we work with $\bar{V}^{\ast}_{k}$ (considering its Skorokhod version, i.e. $\bar{V}^{\ast}_{k}=\bar{Z}^{\ast}_{k}$ a.s. given $Z_{n-k,n}=t$), the definitions of $\overline{o}_{P}$ and $\overline{O}_{P}$ can be easily recast in terms of $\bar{V}^{\ast}_{k}.$
We now provide the extreme value analogue to Theorem 1.1 of [38], where the decomposition is now in terms of sums of functions of upper order statistics, and the statement crucially involves the above-defined concept of uniform convergence in probability.
Theorem 3.1 (Decomposition, truncated case).
Let $\varphi$ have envelope $\bar{\varphi}$ and satisfy Temporary Condition 1. Let $k=k_{n}$ be an intermediate sequence. Then the following random sequence is $\overline{o}_{P}(k^{-1/2})$:
$r_{k,n}=S_{k,n}(\varphi)-\Big{\{}\frac{1}{k}\sum_{i=1}^{k}\varphi(Z_{i,k}^{\ast})\gamma_{0}^{t}(Z_{i,k}^{\ast})\delta^{\ast}_{[i:k]}+\frac{1}{k}\sum_{i=1}^{k}\gamma_{1}^{t}(Z_{i,k}^{\ast})(1-\delta^{\ast}_{[i:k]})-\frac{1}{k}\sum_{i=1}^{k}\gamma_{2}^{t}(Z_{i,k}^{\ast})\Big{\}},$ (3.4)
where $\{\delta^{\ast}_{i}\}_{i=1}^{k}$ are such that their concomitants $\{\delta^{\ast}_{[i:k]}\}_{i=1}^{k}$ equal $\{\delta_{[n-k+i:n]}\}_{i=1}^{k}$ a.s.
We proceed to state a general version of Theorem 3.1 by lifting the truncation condition (3.2). In the proof of this extension, we require analogues of (1.5) and (1.6) in [38], which help to control the tail. For that purpose, we impose the following key condition, which is natural for the limit variance to exist.
@@
$\int_{1}^{\infty}|\bar{\varphi}(x)|^{2+\varepsilon}x^{\alpha(\varepsilon)}\mathop{}\!\mathrm{d}x<\infty,$ (3.5)
@@
Proposition 3.2.
Assume Condition 1. Then for large $t$
$\int_{1}^{\infty}\varphi^{2}(\gamma^{t}_{0})^{2}\mathop{}\!\mathrm{d}H^{1,t}<\infty$ (3.6)
and
$\int_{1}^{\infty}|\varphi|\sqrt{C^{t}}\mathop{}\!\mathrm{d}F^{t}<\infty.$ (3.7)
Theorem 3.3 (Decomposition, general case).
Assume Condition 1, and let $k=k_{n}$ be an intermediate sequence. Then the random sequence $r_{k,n}$ defined by (3.4) is $\overline{o}_{P}(k^{-1/2})$.
The above representation, without additional assumptions allows us to prove that the extreme Kaplan–Meier integral $S_{k,n}(\varphi)$ (3.1) is a consistent estimator for
$S_{\circ}(\varphi):=\int_{1}^{\infty}\varphi\mathop{}\!\mathrm{d}F^{\circ}$

### v5

@@ Abstract
A novel and comprehensive methodology designed to tackle the challenges posed by extreme values in the context of random censorship is introduced. The main focus is on the analysis of integrals based on the product-limit estimator of normalized upper order statistics, called extreme Kaplan–Meier integrals. These integrals allow for the transparent derivation of various important asymptotic distributional properties, offering an alternative approach to conventional plug-in estimation methods. Notably, this methodology demonstrates robustness and wide applicability among various tail regimes. A noteworthy by-product is the extension of generalized Hill-type estimators of extremes to encompass arbitrary tail behavior, which is of independent interest. The theoretical framework is applied to construct novel estimators for real-valued extreme value indices for right-censored data. Simulation studies confirm the asymptotic results and, in a competitor case, mostly show superiority in mean square error. An application to brain cancer data demonstrates that censoring effects are properly accounted for, even when focusing solely on tail classification.
@@
Kaplan–Meier estimator,
@@ 1 Introduction
Analyzing extreme events that experience right censoring is a crucial undertaking in the field of statistics, and failing to account for these censoring effects can lead to significant bias. Roughly speaking, by treating extreme values that remain concealed due to censoring as fully-observed data points, we may underestimate their true effects. Such circumstances commonly arise when investigating weather phenomena, financial markets or re-insurance claim sizes. Likewise, in biomedical applications, the proper classification of the behavior of the largest observations is a delicate task due to incomplete data. Hence, exploring censored extremes is a crucial step in mitigating associated risks. Given the escalating worldwide prevalence of catastrophic occurrences, the study of censored extremes plays an important role in constructing enhanced statistical models for these risks. In turn, these correctly-calibrated models facilitate informed decision making.
@@
The foundational statistical work in fully-observed extreme value theory is by now classic and well-understood in the univariate case; see [2, 14, 11], and to a large degree for univariate time series, as can be found in [24]. Similarly, the pillars of survival analysis have been laid out since the seventies, in different flavors, with the counting process point of view being one of the the most popular approaches, and with martingale techniques playing an ubiquitous role; see [19, 1]. However, a unified methodology for estimating various extreme values under censoring is still lacking. Our paper aims to address this gap.
@@
More precisely, this paper proposes the analysis of randomly right-censored extremes through Kaplan–Meier integrals of normalized upper order statistics (see [33] for the analogous non-extreme case). This generalizes integral functions with respect to the empirical distribution function of normalized upper order statistics, which have received interest in the literature. For instance, such integrals can often be cast as sums of extreme values from iid samples, which were studied in [10] for regularly varying tails and in [9] for arbitrary tail behavior. Upon replacement of the empirical distribution function with the Kaplan–Meier estimator of upper order statistics of a censored sample, the analysis of integral functions is no longer transparent, neither with theory of empirical processes nor using the sums approach. This paper is devoted to studying precisely these integral functions by using the elegant U-statistics approach of [32], together with machinery from extreme value theory.
@@
In order to generalize the approach of [32] to the extreme case, the theory of U-statistics [29] and of residual estimation (based on averages of functions of normalized upper order statistics) of [28] are important. However, they are not directly applicable, the former having to be adapted to accommodate upper order statistics (see [26] for a related concept), and the latter requiring extension to arbitrary tail regimes. These adaptations are provided in this paper, and yield a unifying and transparent treatment of extreme estimators and their convergence, imposing natural conditions. Estimators of the parameter governing the tail behavior (the so-called extreme value index) derived from extreme Kaplan–Meier integrals have similar form and in some cases equal asymptotic variance as the ones treated in [6], though the construction principle is rather different and our theory allows for the direct construction of estimators of further extreme value characteristics.
The central result of the paper is Theorem 3.1, providing a general decomposition of extreme Kaplan–Meier integrals in terms of sums of conditionally i.i.d. random variables. The latter result and its extensions, in combination with other conditions, yield useful limit theorems and applications. The remaining theoretical contributions of the paper can be summarized as follows:
@@
    Suitable conditions are established under which extreme Kaplan–Meier integrals are weakly consistent, Theorem 3.2, and asymptotically normal for distributions with regularly varying tails. Asymptotic normality is provided in two steps, with only the last one, comprised in Theorem 3.6 and Corollary 3.7, requiring a second-order condition. The main unavoidable restriction is a second moment condition for the limit variance to exist. As an illustration, the asymptotic normality of a Hill-type estimator initially proposed by [35] for the positive extreme value index is established for the first time in Corollary 3.9. Another notable application is Corollary 3.11 where the conditions of consistency and asymptotic normality of the extreme Kaplan–Meier estimator are provided.
@@
    Residual estimation theory is generalized to arbitrary tail regimes and in particular it is found that the convergence rates vary across regimes. The results are presented for fully-observed data and are of independent interest. Here, the main result is Theorem 4.4. The exposition is in terms of so-called candidate functions, which themselves must satisfy a second-order type condition. This extension is key, since residual estimators are the uncensored version of extreme Kaplan–Meier integrals.
@@
    The results of Section 3 are extended to all max-domains of attraction in terms of candidate functions, using generalized residual estimators, now allowing for censored data. Here, certain regularly varying functions with tail index that can be identified as the asymptotic non-censoring proportion play a central role; see Condition 4. We highlight Theorem 5.2 and Theorem 5.3 as our main extensions.
@@
    Finally, we provide a larger application, by establishing full asymptotics for a novel censored version of the moment estimator of [12], using all the main components of our theory. We provide in Theorem 6.1 consistency and in Theorem 6.2 asymptotic normality.
The cross-fertilization of extreme value theory and survival analysis has been explored before in various contexts. We provide an account of some of the most important contributions. In [3] the Pareto index estimation was considered for the first time under random right-censorship framework (see also [4, 13] covering arbitrary tail regimes), which has since produced some generalizations, e.g. using covariates in [25], truncation in [20], trimming in [8], and bias-reduction in [5, 21]. Another approach, which neatly formulates the censoring problem with extremes as a path-wise problem, was proposed in [35] through constructions based on expected values, while [36] developed a competing risks model using techniques from [33]. Subsequently, a partial result on the asymptotic distribution was recently provided in [6]. In terms of missing extreme data, the main reference is [37], and finally the machinery of extreme value theory has recently been imported to survival analysis in a different context, namely for cure-rate modelling in the case of insufficient follow-up; see [15, 16, 17].
The focus of the existing literature is on providing methods for estimating the extreme value index in censored samples, or estimators derived thereof. In contrast, our results facilitate further directions (apart from the presented Hill-type and moment estimators) such as: the estimation of the scale and second-order parameters; the construction of estimates for distribution tails; estimation of extreme quantiles and other extreme characteristics for censored samples. Moreover, popular methods in extreme value theory such as maximum likelihood estimation and probability weighted moments can be extended to censored extremes using our methodology.
Organisation of the paper. In Section 2 we introduce notation and results from extreme value theory and survival analysis relevant to our setup. Section 3 is devoted to the study of extreme Kaplan–Meier integrals when both event and censoring variables have distributions with regularly varying tails. Residual estimators, the uncensored versions of extreme Kaplan–Meier integrals, are extended in Section 4 to arbitrary tail regimes, and their applications are found in the Supplementary Material. Extensions of the results of Section 3 are finally provided by generalizing Section 4 for censored data; this is done in Section 5. Finally, Section 6 provides a prime application of our theoretical results, proposing and analysing in detail an extreme value index estimator for censored data and for all max-domains of attraction. Three simulation studies are provided in Section 7, and an application to a brain cancer dataset is given in Section 8. All proofs may be found in the Supplementary Material.
@@ 2 Preliminaries and setting
This section provides standard preliminaries from survival analysis and extreme value theory, and may be used as a reference for notation in the remainder of the paper.
@@ 2.1 Random censoring
We assume the random right-censoring setting, that is, we observe the following datapoints
@@
where $\{X_{i}\}_{i=1}^{n}$ and $\{Y_{i}\}_{i=1}^{n}$ are independent i.i.d. sequences cdf $F$ and $G$, respectively. Notice that $\delta_{i}$ indicates whether the variable of interest $X_{i}$ has been observed or not. Let $H$ denote the cdf of $Z_{1}$, which by independence of the variable $X_{1}$ and the censoring mechanism $Y_{1},$ is given by
@@
A consistent estimator of $F$, [34], is now provided by the Kaplan–Meier product-limit estimator $\amsmathbb{F}_{n}$ defined by
@@
Let $\varphi:\amsmathbb{R}\to\amsmathbb{R}$ be some measurable function. Write $F\{a\}=F(a)-F(a-)$, such that $F\{a\}>0$ iff $a$ is an atom of $F$. Let $A$ be the set of all atoms of $H,$ possibly empty. Under $\int|\varphi|\mathop{}\!\mathrm{d}F<\infty,$ Stute and Wang [33] showed that
@@
where $(Z,\delta)\stackrel{{\scriptstyle d}}{{=}}(Z_{1},\delta_{1}).$ The following three functions are necessary to state a key asymptotic decomposition of $\int\varphi\,\mathop{}\!\mathrm{d}\amsmathbb{F}_{n}$, and are provided here for future reference,
@@ 2.2 Extreme value theory
When analyzing extremes, a popular approach is to consider a categorization in terms of the possible non-degenerate limits of the excess distribution. It turns out that there are only three such limits, which can be written in terms of the Generalised Pareto Distribution (GPD), with cumulative distribution function given by
$G_{\gamma}(y)=\begin{cases}1-\left(1+\gamma\frac{y-\mu}{\sigma}\right)^{-1/\gamma},&\text{if }\gamma\neq 0,\\ 1-\exp\left(-\frac{y-\mu}{\sigma}\right),&\text{if }\gamma=0,\end{cases}$ (2.9)
for $y\geq\mu$ when $\gamma\geq 0$, and $\mu\leq y\leq\mu-\sigma/\gamma$ when $\gamma<0$, where $\mu\in\amsmathbb{R}$ is the location parameter, $\sigma>0$ is the scale parameter, and $\gamma\in\amsmathbb{R}$ is the shape parameter called the extreme value index.
The following celebrated result then links the excesses over a large threshold with the GPD distributions.
Theorem 2.1 (Pickands–Balkema–de Haan Theorem).
Let $X$ be a random variable with distribution function $F$. Then the only possible non-degenerate limit of $\amsmathbb{P}(X-u\leq a(u)x+b(u)\mid X>u)$ with measurable functions $a(u)>0$ and $b(u)$ is $G_{\gamma}$.
The parameter $\gamma$ not only governs the tail behavior, but also categorizes distribution into three distinct so-called max-domains of attraction (MDA). Thus, we say that $F$ belongs to the Fréchet, Gumbel, or Weibull MDA if $\gamma>0$, $\gamma=0$, or $\gamma<0$, respectively. We then write $F\in\mathcal{D}(G_{\gamma})$.
Similarly, for any cdf $F$ define its tail counterpart as
@@
which is itself a cdf and of practical utility in the sequel. Further, let $\gamma^{t}_{0},$ $\gamma^{t}_{1}$ and $\gamma^{t}_{2}$ be defined similarly to (2.4), (2.5) and (2.6), by replacing $H,H^{0}$ and $H^{1}$ with $H^{t},$ $H^{0,t}$ and $H^{1,t},$ respectively, where
@@
3 Extreme Kaplan–Meier Integrals
This section studies the consistency and asymptotic normality of extreme Kaplan–Meier integrals for regularly varying tails, with the main results being the decomposition of Theorem 3.1, and the asymptotic normality of Corollary 3.7, though the intermediate results are of independent interest.
We consider the case where the variables $X_{i}$ and the censoring mechanism $Y_{i}$ are both heavy-tailed, i.e. belonging to the Fréchet max-domain of attraction. In later sections this restriction is relaxed to also allow them to belong to the Gumbel and Weibull max-domains of attraction. More specifically, we assume now that $F,$ $G$, and hence $H$, are regularly varying cdfs with respective tail indices $1/\gamma_{F},$ $1/\gamma_{G}$ and $1/\gamma_{H},$ where by (2.1) we obtain
@@ 3 Extreme Kaplan–Meier Integrals
In this case, by regular variation of $F,$
$1-F^{t}(x)\to x^{-1/\gamma_{F}}\quad\text{ as }\quad t\to\infty,$ (3.1)
and similarly for $G$ and $H$.
Let $Z_{1,n}\leq\ldots\leq Z_{n,n}$ be the order statistics of a randomly right censored sample $\{(Z_{i},\delta_{i})\}$ and let $\{\delta_{[i:n]}\}_{i=1}^{n}$ be the concomitants of $\{Z_{i,n}\}_{i=1}^{n}.$ We are interested in analyzing the extreme value behavior of the datapoints $X_{1},\dots,X_{n}$, which is possible through the study of $Z_{i,n}$ and $\delta_{[i:n]}$. The following estimator targets the tail of the distribution while accounting for censoring effects.
@@
The extreme Kaplan–Meier estimator (EKM) is given by [^1] We assume $10^{0}=$ 1 by convention.
@@
Let $F,G,$ and therefore $H$ be eventually continuous (i.e. continuous to the right of a large enough value). The asymptotic behavior of extreme Kaplan–Meier integrals is studied below, defined as integrals with respect to the EKM estimator:
@@
It is easily seen from (3.2) that
$S_{k,n}(\varphi)=\sum_{i=1}^{k}\omega_{ik}\;\varphi\left(\frac{Z_{n-i+1,n}}{Z_{n-k,n}}\right),$ (3.4)
where for $1\leq i\leq k,$
$\omega_{ik}=\frac{\delta_{[n-i+1:n]}}{i}\prod_{j=i+1}^{k}\left[\frac{j-1}{j}\right]^{\delta_{[n-j+1:n]}}.$ (3.5)
We assume that $\varphi$ is a continuous and eventually monotone function on $[1,\infty).$
We now provide the extreme value analogue to Theorem 1.1 of [32], where the decomposition is now in terms of sums of functions of upper order statistics. Setting $\varphi=\log$ in (3.3), we get a version of the celebrated Hill estimator [22] of positive $\gamma_{F}$ in presence of censoring, see Corollary 3.9 for details and also [13] and [35] for related estimators.
We impose the following key condition, which is natural for the limit variance to exist.
@@
$\int_{1}^{\infty}(\varphi(x))^{2}\,x^{\alpha(\varepsilon)}\mathop{}\!\mathrm{d}x<\infty,$ (3.6)
@@
We call the sequence of integers $k_{n}$ intermediate if
Theorem 3.1 (Key decomposition).
Let $F$ and $G$ be regularly varying and eventually continuous cdfs, and $\varphi$ be continuous and eventually monotone. Assume Condition 1, and let $k=k_{n}$ be an intermediate sequence. Then the following random sequence
$r_{k,n}$ $=$ $S_{k,n}(\varphi)$ (3.7)
$-\Big{\{}\frac{1}{k}\sum_{i=1}^{k}\varphi(Z_{i,k}^{\ast})\gamma_{0}^{t}(Z_{i,k}^{\ast})\delta_{[n-k+i:k]}\!+\!\frac{1}{k}\sum_{i=1}^{k}\gamma_{1}^{t}(Z_{i,k}^{\ast})(1\!-\!\delta_{[n-k+i:k]})\!-\!\frac{1}{k}\sum_{i=1}^{k}\gamma_{2}^{t}(Z_{i,k}^{\ast})\Big{\}}$
is such that for every $\epsilon>0$
$\lim_{n\to\infty}\limsup\limits_{t\rightarrow\infty}{\amsmathbb{P}}\big{(}|k^{1/2}r_{k,n}|>\epsilon|\,Z_{n-k,n}=t\big{)}=0.$
Here, $(Z^{\ast}_{1,k},\dots,Z^{\ast}_{k,k})=(Z_{n-k+1,n}/Z_{n-k,n},\ldots,Z_{n,n}/Z_{n-k,n})$.
In simple words, Theorem 3.1 states that $S_{k,n}(\varphi)$ given $Z_{n-k,n}=t$ can be represented as a sum of i.i.d. random variables plus a remainder term, which is negligible of order $k^{-1/2}$ for large $n$, uniformly in large $t$ (confer with the proof for further details). The above representation, without additional assumptions, allows us to prove that the extreme Kaplan–Meier integral $S_{k,n}(\varphi)$ (3.3) is a consistent estimator for
$S_{\circ}(\varphi):=\int_{1}^{\infty}\varphi\mathop{}\!\mathrm{d}F^{\circ},$